OpenStudy (anonymous):

One jet left Detroit at 6:30 a.m. flying south at 410 mph, and another left at 8:30 a.m. flying south at 620 mph. How many hours will it take the second jet to catch up to the first? Round to the nearest tenth if necessary.

hartnn (hartnn):

u know the formula speed=distance/time here distances traveled are equal so equate the product of speed and time for both....

OpenStudy (anonymous):

how do i equate the distance?

hartnn (hartnn):

sorry, u have the speed,u just need time for both planes now... then u can equate product of spped and time for both planes...

OpenStudy (anonymous):

whats the answer then?

OpenStudy (anonymous):

r t = d Let t be the time the second jet takes to overtake the first. Then: 410 (t + 2) = 620 t t = 82/21=3.90476 or 3.9 hours to the nearest tenth

OpenStudy (anonymous):

Thank you for the medal, @reonrobe.